OpenStudy (anonymous):

Is (3,6) a solution for y < 2/3x + 4? Please help I don't get it

OpenStudy (tkhunny):

Simply substitute the given values and see if it works. y < (2/3)x + 4 Substitute x = 3 y < (2/3)(3) + 4 = 2 + 4 = 6 y < 6 Substitute y = 6 6 < 6 -- Well, that doesn't make sense!

OpenStudy (anonymous):

Thank You!!